# Procountable groups are not classifiable by countable structures

### v2

@@ 2.4. Further remarks
Such equivalence relations were called complete analytic equivalence relations in [5]. Here we call them universal to avoid possible confusion. In [5] and subsequent research, many natural equivalence relations were found to be universal analytic equivalence relations. In particular, it was shown in [5] that the uniform homeomorphism for all Polish metric spaces is a universal analytic equivalence relation.
@@
Toward a contradiction, assume that there is a map $f$ witnessing that $c_{0}$ is Borel reducible to $\mbox{\rm GI}\times\ell_{\infty}$. The map $f$ induces a homomorphism of equivalence relations from $c_{0}$ to GI. By the Hjorth turbulence theorem (see [13, Theorem 3.18 and Example 3.23]; also see [8, Theorems 10.4.2 and 10.5.2]), there is a $c_{0}$-invariant comeager set $C$ on which $c_{0}$ is sent by $f$ to a single GI-class. This implies that $c_{0}$ on $C$ is Borel reducible to $\ell_{\infty}$. However, $c_{0}\upharpoonright C^{2}$ is ${\bf\Pi}^{0}_{3}$-complete (see [8, Exercise 8.5.1]) and as such, it is not Borel reducible to any $F_{\sigma}$ equivalence relation (see [8, Theorem 8.5.2] and its proof).
@@ 3.2. Proω categories
We provide some preliminaries on the correspondence of isomorphisms of procountable groups and certain isomorphisms of inverse systems of type $\omega$. Our purpose is mainly to lay down a notational framework; more general results are obtained in [21] using a somewhat different language.
@@
commute. Composition of pre-morphisms is defined in the canonical way (see [21, pg. 5]), and the identity pre-morphism is $({\mathrm{id}_{\overline{A}}},\mathrm{id}_{\mathbb{N}})$, where $\mathrm{id}_{\overline{A}}=\{{\mathrm{id}_{A_{n}}}\}_{n\in\mathbb{N}}$.
@@
The transitivity of the equivalence relation $\sim$ and its compatibility with composition of pre-morphisms are proved in [21, pg. 6]. The morphisms of $\mathrm{Pro}_{\omega}(\mathcal{K})$ are defined as the equivalence classes of pre-morphisms under $\sim$.
@@
in $\mathrm{Pro}_{\omega}(\mathrm{GrpEpi})$. Then the composition $(\bar{g},\psi)\circ(\bar{f},\phi)$, which we recall to equal
@@ 4.1. The 2-nilpotent groups $L(X)$
As noted in Mekler [22], the centre $Z(N)$ of $N$ is an elementary abelian $p$-group (and hence an $\mathbb{F}_{p}$ vector space) with basis $x_{r,s}$ for $r<s$. Given an undirected irreflexive graph $A$ with vertex set $\mathbb{N}$, Mekler sets
@@
The following fact already noted in [22] is evident.
@@
We will only use Mekler’s construction for a very simple infinite graph: the one such that each connected component has size 2. (While we will no use it, we note that this graph is “nice” in the sense of Mekler, cf. [22, Definition 1.1].)
@@
The idea is that $G(A)$ has the minimum commutativity: any pair of elements that commute nontrivially have to generate the same subgroup as some pair $x_{v},x_{v+1}$ for even $v$, all modulo the centre.
@@
We may assume that $c=\prod_{r\in C}{x_{r}}^{{\alpha}_{r}}$ and $d=\prod_{s\in D}{x_{s}}^{{\alpha}_{s}}$ as in 4.1(ii). If $r\in\mathbb{N}\setminus C$ we let $\alpha_{r}=0$, and if $s\in\mathbb{N}\setminus D$ we let $\beta_{s}=0$. Suppose the first condition fails. We show that $\underline{c}=\underline{d}^{\lambda}$ for some $\lambda\in\mathbb{N}$.
By assumption’ there is no $k\in\mathbb{N}$ such that both $C$ and $D$ are contained in $\{2k,2k+1\}$. Note that neither $C$ nor $D$ can be a singleton: if (say) $C=\{r\}$, then there is $s\in D$ such that $r^{*}\neq s^{*}$. Then $\alpha_{r}\beta_{s}\neq 0$ while $\alpha_{s}\beta_{r}=0$, which contradicts the assumption that $c,d]\in Z(L)$ because of 4.2.
@@
Given a countably infinite set $X$, let $N(X)$ be the free 2-nilpotent group of exponent $p$ on distinct free generators $a_{v},b_{v}$ ($v\in X$). Let
@@
Let $\mathcal{L}$ be the category of groups of the form $L(X)$ with morphisms the epimorphism $g\colon L(X)\to L(Y)$ such that for each $v\in X$, the rank of $\langle\underline{g(a_{v})},\underline{g(b_{v})}\rangle$ as a subgroup of $L(Y)/Z(L(Y))$ equals $2$. A functor $F$ from $\mathcal{L}$ to the category of countable sets with surjections is declared by sending $L(X)$ to $X$, and sending a morphism $g:L(X)\to L(Y)$ to the map $p=F(g)\colon X\to Y$ such that
@@
To show that $F$ is well-defined for morphisms of $\mathcal{L}$, it suffices to apply 4.4 to $L=L(Y)$, $c=g(a_{v}),d=g(b_{v})$, noting that $[c,d]=e$. One verifies easily that $F$ is compatible with composition, and that $F(\widehat{q})=q$. ∎
@@ 4.2. The reduction for the second step
Let $r_{n}\colon T_{\phi(n)}\to U_{n}$ and $s_{n}\colon U_{\psi(n)}\to T_{n}$ be defined by $r_{n}=F(f_{n})$ and $s_{n}=F(g_{n})$ where $F$ is the functor from Lemma 4.6. By that lemma we obtain commutative diagrams of surjections between sets
@@
The $\Phi$ and $\Psi$ are uniformly continuous by 2.5. By ($\star$4.2) and 4.6 we have $p_{\phi(\psi(n)),n}=s_{n}\circ r_{\psi(n)}$ and $q_{\psi(\phi(n)),n}=r_{n}\circ s_{\phi(n)}$ for each $n$. Hence $\Psi\circ\Phi$ is the identity on $[T]$, and $\Phi\circ\Psi$ is the identity on $[U]$. So $T]\cong_{u}[U]$ as required. ∎

### v3

@@ 2.4. Further remarks
Such equivalence relations were called complete analytic equivalence relations in FLR . Here we call them universal to avoid possible confusion. In FLR and subsequent research, many natural equivalence relations were found to be universal analytic equivalence relations. In particular, it was shown in FLR that the uniform homeomorphism for all Polish metric spaces is a universal analytic equivalence relation.
@@
Toward a contradiction, assume that there is a map $f$ witnessing that $c_{0}$ is Borel reducible to $\mbox{\rm GI}\times\ell_{\infty}$. The map $f$ induces a homomorphism of equivalence relations from $c_{0}$ to GI. By the Hjorth turbulence theorem (see [Hj, , Theorem 3.18 and Example 3.23]; also see [GaoBook, , Theorems 10.4.2 and 10.5.2]), there is a $c_{0}$-invariant comeager set $C$ on which $c_{0}$ is sent by $f$ to a single GI-class. This implies that $c_{0}$ on $C$ is Borel reducible to $\ell_{\infty}$. However, $c_{0}\upharpoonright C^{2}$ is ${\bf\Pi}^{0}_{3}$-complete (see [GaoBook, , Exercise 8.5.1]) and as such, it is not Borel reducible to any $F_{\sigma}$ equivalence relation (see [GaoBook, , Theorem 8.5.2] and its proof).
@@ 3.2. Proω categories
We provide some preliminaries on the correspondence of isomorphisms of procountable groups and certain isomorphisms of inverse systems of type $\omega$. Our purpose is mainly to lay down a notational framework; more general results are obtained in Mardesic.Segal using a somewhat different language.
@@
commute. Composition of pre-morphisms is defined in the canonical way (see [Mardesic.Segal, , pg. 5]), and the identity pre-morphism is $({\mathrm{id}_{\overline{A}}},\mathrm{id}_{\mathbb{N}})$, where $\mathrm{id}_{\overline{A}}=\{{\mathrm{id}_{A_{n}}}\}_{n\in\mathbb{N}}$.
@@
The transitivity of the equivalence relation $\sim$ and its compatibility with composition of pre-morphisms are proved in [Mardesic.Segal, , pg. 6]. The morphisms of $\mathrm{Pro}_{\omega}(\mathcal{K})$ are defined as the equivalence classes of pre-morphisms under $\sim$.
@@
in $\mathrm{Pro}_{\omega}(\mathrm{GrpEpi})$. Then the composition $(\bar{g},\psi)\circ(\bar{f},\phi)$, which we recall equals
@@ 4.1. The 2-nilpotent groups $L(X)$
As noted in Mekler Mekler1985 , the centre $Z(N)$ of $N$ is an elementary abelian $p$-group (and hence an $\mathbb{F}_{p}$ vector space) with basis $x_{r,s}$ for $r<s$. Given an undirected irreflexive graph $A$ with vertex set $\mathbb{N}$, Mekler sets
@@
The following fact already noted in Mekler1985 is evident.
@@
We will only use Mekler’s construction for a very simple infinite graph: the one such that each connected component has size 2. (While we will no use it, we note that this graph is “nice” in the sense of Mekler, cf. [Mekler1985, , Definition 1.1].)
@@
The idea is that $G(A)$ has the minimum commutativity: any pair of elements that commute nontrivially have to generate the same subgroup as some pair $x_{i},x_{i+1}$ for even $i$, in the central quotient.
@@
We may assume that $c=\prod_{r\in C}{x_{r}}^{{\alpha}_{r}}$ and $d=\prod_{s\in D}{x_{s}}^{{\alpha}_{s}}$ as in 4.1(ii). If $r\in\mathbb{N}\setminus C$ we let $\alpha_{r}=0$, and if $s\in\mathbb{N}\setminus D$ we let $\beta_{s}=0$. Suppose there is no even $i\in\mathbb{N}$ such that $\langle\underline{c},\underline{d}\rangle=\langle\underline{x}_{i},{\underline{x}}_{i+1}\rangle$. We show that $\underline{c}=\underline{d}^{\lambda}$ for some $\lambda\in\mathbb{N}$.
By assumption there is no $k\in\mathbb{N}$ such that both $C$ and $D$ are contained in $\{2k,2k+1\}$. Note that neither $C$ nor $D$ can be a singleton: if (say) $C=\{r\}$, then there is $s\in D$ such that $r^{*}\neq s^{*}$. Then $\alpha_{r}\beta_{s}\neq 0$ while $\alpha_{s}\beta_{r}=0$, which contradicts the assumption that $c,d]\in Z(L)$ because of 4.2.
@@
Given a countably infinite set $X$, let $N(X)$ be the free 2-nilpotent group of exponent $p$ on distinct free generators $\{a_{v},b_{v}\colon v\in X\}$. Let
@@
Let $\mathcal{L}$ be the category of groups of the form $L(X)$ with morphisms the epimomorphisms $g\colon L(X)\to L(Y)$ such that for each $v\in X$, the rank of $\langle\underline{g(a_{v})},\underline{g(b_{v})}\rangle$ as a subgroup of $L(Y)/Z(L(Y))$ equals $2$. A functor $F$ from $\mathcal{L}$ to the category of countable sets with mappings between them is declared by sending $L(X)$ to $X$, and sending a morphism $g:L(X)\to L(Y)$ to the map $p=F(g)\colon X\to Y$ such that
@@
To show that $F$ is well-defined for a morphism $g$ of $\mathcal{L}$ as above, it suffices to apply 4.4 to $L=L(Y)$, $c=g(a_{v}),d=g(b_{v})$, noting that $[c,d]=e$. Thus in $L(Y)/Z(L(Y))$ we have $\langle\underline{c},\underline{d}\rangle=\langle\underline{a}_{y},\underline{b}_{y}\rangle$ for a unique $y\in Y$, and hence $p(v)=y$ is defined. One verifies easily that $F$ is a functor, and that $F(\widehat{q})=q$. ∎
@@ 4.2. The reduction for the second step
Let $r_{n}\colon T_{\phi(n)}\to U_{n}$ and $s_{n}\colon U_{\psi(n)}\to T_{n}$ be defined by $r_{n}=F(f_{n})$ and $s_{n}=F(g_{n})$ where $F$ is the functor from Lemma 4.6. By that lemma we obtain commutative diagrams of maps between sets
@@
Then $\Phi$ and $\Psi$ are uniformly continuous by 2.5. By ($\star$4.2) and 4.6 we have $p_{\phi(\psi(n)),n}=s_{n}\circ r_{\psi(n)}$ and $q_{\psi(\phi(n)),n}=r_{n}\circ s_{\phi(n)}$ for each $n$. Hence $\Psi\circ\Phi$ is the identity on $[T]$, and $\Phi\circ\Psi$ is the identity on $[U]$. So $T]\cong_{u}[U]$ as required. ∎